if Linda takes 10 minutes to jog 1 mile, what is her speed in miles per hour?
divide 4/5 by 2/3?
Subtract 5 2/3 from 8/ 1/4
what is the LCM of 12 and 18?
for a class picnic, the following wraps were made, 8 ham, 8 veggie, 15 turkey, and 19 tuna. What percent of wraps are turkey?

Answer by stanbon(75887)

if Linda takes 10 minutes to jog 1 mile, what is her speed in miles per hour?
-----
(1 mile/10 min)(60 min/1 hr) = (60/10) mph = 6 mph
----------------------------
divide 4/5 by 2/3?
(4/5)/(2/3) = (4/5)*(3/2) = 12/10 = 6/5 = 1 1/5
----------------------------
Subtract 5 2/3 from 8/ 1/4
8 1/4 - 5 2/3 = (33/4)-(17/4) = 16/4 = 4
----------------------------
what is the LCM of 12 and 18?
12 = 2^2*3
18 = 2*3^2
LCM = 2^2*3^2 = 4*9 = 36
----------------------------
for a class picnic, the following wraps were made, 8 ham, 8 veggie, 15 turkey, and 19 tuna. What percent of wraps are turkey?
Total = 8+8+15+19 = 50
Percept turkey = 15/50 = 30/100 = 30%
============================
Cheers,
Stan H.